what is the addition method if possible?
3x+9y=9
-x+5y=-3

Answer by dondon(1)

3x + 9y = 9 (number this equation as equation i)
-x + 5y = -3 (number this equation as equation ii)
Now multiply equation ii) by 3 which gives:
-3x + 15y = -9 (number this equation as equation iii)
Now add equation i) and equation iii) which gives:
24y = 0
=> y = 0
Now plug this obtained value of y in any of equation i) or ii) to get value of x:
plugging in i)
3x + 9(0) = 9
=> 3x = 9
=> x = 3
so we get x = 3 & y = 0 as the answer.